OpenStudy (anonymous):

i need help with this equation CAN SOMEBODY HELP ME OUT: 2x 0VER 5x^2-37x+14 minus 3x OVER5x^2-33x-14?? in simplest form

OpenStudy (anonymous):

\[\frac{2 x}{(x-7) (5 x-2)}-\frac{3 x}{(x-7) (5 x+2)} \]The above can be factored.\[\frac{x}{x-7}\left(\frac{2}{5x-2}-\frac{3}{5 x+2}\right) \]\[\frac{x}{x-7}\left(-\frac{5 (x-2)}{(5 x-2) (5 x+2)}\right) \]\[-\frac{5x (x-2) }{(x-7) (5 x-2) (5 x+2)} \]